OpenStudy (anonymous):

store A, store B, and store C sell soccer balls. the ratio of the number of soccer balls sold of A to B is 7/4. The ratio of the number of soccer balls sold of B to C is 5/3. if store C sold 36 soccer balls, how many soccer balls did store A sell?

OpenStudy (anonymous):

nice question: a/b=7/4 b/c = 5/3 c=36 b=36/3*5=60 a = 60/4 *7 =15*7=105

OpenStudy (anonymous):

thank u so much...:D